One year ago, Alpha Supply issued 15-year bonds at par. The bonds have a coupon rate of 6.5 percent, paid semiannually, and a face value of $1,000. Today, the market yield on these bonds is 7.2 percent. What is the percentage change in the bond price over the past year?

5.94 percent

5.38 percent

-6.11 percent

-5.87 percent

The bond price did not change.
Answer #1

Solution:

One year ago the bond price is:

C=6.5%×1000=65/2=32.5 semiannual

r=7.2/2=3.6%

n=30-2=28

PV=C[(1-(1+r)^-n)/r+FV(1+r)^-n

= $32.5[(1-(1+.036)^-28)/.036+1000(1+.036)^-28

= $32.5(17.4591) +1000(.37147)

= $567.42 +$371.47

=$938.89

%change in bond price=$(938.89-1000)/1000

=-6.11%